OpenStudy (anonymous):

(-6.2+5.9i)-(2.4+9.1i)

OpenStudy (anonymous):

have to subtract the complex expreesion and simplify which i got -11.8

OpenStudy (anonymous):

same thing dude = -6.2 + 5.9i - 2.4 - 9.1i = -8.6 - 32i

OpenStudy (anonymous):

its like same old algerbraic expression.. so chill.. combine imaginary terms and constant terms separatetly